OpenStudy (anonymous):

In rectangle TUVW, (angle)TXU = 100(degrees) and TV =14inches. Find the measures below.

OpenStudy (alexandervonhumboldt2):

hmm drwing?

OpenStudy (alexandervonhumboldt2):

what to find?

OpenStudy (anonymous):

A. TX= B. (angle)TXW= C. WX=

OpenStudy (alexandervonhumboldt2):

TX=XU=WX=XV=1/2TV=1/2*14=7 A. TX=1/2*14=7 B. (angle)TXW=180-100=80 C. WX=1/2TV=7

OpenStudy (mathstudent55):

The diagonals of a rectangle bisect each other.

OpenStudy (alexandervonhumboldt2):

yes for 2 use defenition of supplementary angles

OpenStudy (mathstudent55):

Also, the diagonals of a rectangle are congruent.